OpenStudy (anonymous):

1/4x - 8 + 2/4x = 19

OpenStudy (anonymous):

x=36

OpenStudy (anonymous):

is tht right

OpenStudy (ahsome):

Equation:\[\frac{1}{4x}-8+\frac{2}{4x}=19\] Convert all to have \(4x\) as the denominator\[\frac{1}{4x}+\frac{-8}{1}*\frac{4x}{4x}+\frac{2}{4x}=\frac{19}{1}*\frac{4x}{4x}\] \[\frac{1}{4x}+\frac{-32x}{4x}+\frac{2}{4x}=\frac{76x}{4x}\] Cancel the denominators: \[1-32x+2=76x\] Move the \(-32x\) to the RHS \[1+2=108x\] Simplify, then solve \[3=108x\] \[x=\frac{3}{108}\]

OpenStudy (ahsome):

Nope, I must have got something wrong then. NVM

OpenStudy (anonymous):

ok

OpenStudy (ahsome):

Is the equation: \[\frac{1}{4x}-8+\frac{2}{4x}=19\]?

OpenStudy (ahsome):

If so, \[x=\frac{1}{36}\] GTG, hopefully someone else can help you. Sorry :(

OpenStudy (anonymous):

ok